The width of a rectangle is 2/3 times the length. The total perimeter of the rectangle is 252cm. What is the width of the rectangle?
Answer by Cromlix(4381)

Hi there,
Make length = x
width = 2/3x
Perimeter = 2(width) + 2(length)
252 = 2(2/3x) + 2(x)
252 = 4/3x + 2x
252 = 4/3x + 6/3x
252 = 10/3x
x = (252 * 3)/10 (* times)
x = 75.6cm
Length = 75.6cm
Width = 50.4cm
Hope this helps :-)